I want to build a model teepee for a school project. I plan to use lengths of flexible tubing for the supports on the teepee and I want the teepee to be 28 inches across and 28 inches tall. How long should the pieces of flexible tubing be?
Answer by nerdybill(7384)

I want to build a model teepee for a school project. I plan to use lengths of flexible tubing for the supports on the teepee and I want the teepee to be 28 inches across and 28 inches tall. How long should the pieces of flexible tubing be?
Apply Pythagorean theorem:
Let x = length of flexible tubing
then
x^2 = 14^2 + 28^2
x^2 = 196 + 784
x^2 = 980
x = 31.3 inches